# Migration Pathways as Diagnostic Probes: Benchmarking and Data-Efficient Training of Specialist vs. Generalist Machine-Learned Force Fields

### v1

@@ Abstract
keywords:
Machine-learned force fields (MLFFs), Foundation models, Fine-tuning, Migration pathways, 2D materials
@@ 1 Introduction
Machine-learned force fields (MLFFs)[10, 11, 12, 13, 14] offer a path to overcome this limitation, bringing near-ab initio accuracy to molecular dynamics simulations at reduced cost. The advent of large pre-trained “foundation models” such as CHGNet[15], MACE-MP[11], MatterSim[16], and UMA[17] marks a paradigm shift: instead of building narrow, system-specific models, one may transfer broad chemical knowledge to a material of interest through fine-tuning. Yet, the choice between building a specialist MLFF from scratch and adapting a generalist foundation model remains unresolved. Questions of data efficiency, stability, and transferability—particularly for non-equilibrium processes—are still open.
Existing benchmarks such as Matbench[18] evaluate global accuracy but do not directly probe the kinetic and dynamical processes most relevant to materials applications. Moreover, MLFF failures are often diagnosed post hoc, when molecular dynamics trajectories diverge unphysically, leading to a brute-force cycle of retraining with ever-larger datasets. This is neither efficient nor informative, as it provides little guidance on where and why the model failed.
@@ 2.1 Data Generation and Model Training
Reference data were generated using Density Functional Theory (DFT) with the PBE functional, comprising  20,000 configurations from ab initio molecular dynamics (AIMD) simulations at multiple temperatures (300K, 600K, 1200K). We employed the MACE architecture for all models due to its state-of-the-art accuracy in representing complex atomic interactions.
@@
    FT-MultiT: Foundation model fine-tuned on multi-temperature data
@@ 2.2 Evaluation Framework
Models were assessed through: (i) molecular dynamics simulations at 300K and 600K to evaluate structural stability and transport properties, (ii) nudged elastic band calculations for migration barriers, and (iii) representation learning analysis using t-SNE visualization of physically interpretable descriptors. Performance metrics included force/energy accuracy, dynamic stability, barrier prediction errors, and latent space clustering quality. Detailed computational parameters are provided in Appendix B.
@@ 3 Results and Discussions
3.1 Benchmarking the MLFF Performance in Equilibrium MD Simulations
To make systematic evaluations benchmarking the model performance, we first assessed the performance of each training strategy in molecular dynamics (MD) simulations (Fig. 1a). This provides a comprehensive test of a model’s ability to generate a stable thermodynamic ensemble and accurately reproduce key structural, dynamic, and transport properties. Each model was used to drive a 200 ps simulation, and the resulting trajectories were analyzed as shown in Fig. 1.
Figure 1: Comprehensive Benchmarking of MLFF Performance in Molecular Dynamics Simulations. (a) A schematic of the unified evaluation protocol. Four candidate models—a zero-shot foundation model, a bespoke model trained from scratch, and two fine-tuned variants—are each used to drive a 200 ps molecular dynamics simulation. The resulting trajectories are then subjected to a uniform set of post-processing analyses to evaluate key physical properties. (b) Thermal transport properties, showing the Heat Flux Autocorrelation Function (HFACF) and its running integral to compute thermal conductivity ($\kappa$). (c) Thermodynamic stability, demonstrated by the evolution of temperature and pressure over the simulation, which remain stable around their target values for all models. (d) The Total Radial Distribution Function (RDF), $g(r)$, revealing the local atomic structure. (e) The Mean Squared Displacement (MSD), used to assess atomic mobility and calculate the diffusion coefficient. (f) The Velocity Autocorrelation Function (VACF), which describes the system’s underlying dynamics.
All MACE models—foundation, scratch, and fine-tuned variants—successfully reproduce the equilibrium structure of Cr-doped Sb2Te3, with RDF analysis showing excellent agreement with AIMD references for all atomic pair correlations (see Appendix A for detailed analysis). While all models maintain stable thermodynamic ensembles at 600K (Fig. 1c), the zero-shot foundation model exhibits a persistent pressure offset, likely due to its training on 0K equilibrium structures rather than finite-temperature configurations.
Despite similar performance on local structural (RDF) and short-time dynamic (VACF) properties, the models diverge significantly in their predictions of long-timescale transport phenomena. Fine-tuned models predict higher diffusion coefficients than both foundation and scratch models (Fig. 1e), with the multi-temperature fine-tuned variant showing the largest enhancement. This suggests that exposure to high-temperature training data creates a flatter potential energy (PES) that persists even at lower temperatures.
@@ 3.1 Benchmarking the MLFF Performance in Equilibrium MD Simulations
3.2 Benchmarking the MLFF for Non-Equilibrium Diffusion Pathways
@@
3.3 Quantitative Comparison of Migration Barriers
@@
3.3.1 The Critical Role of Task-Specific Fine-Tuning
@@
3.3.2 Further Testing Different Trajectories
To probe the generalization capabilities of the different MLFF training strategies, we evaluated their performance on two distinct Cr migration pathways (Fig. LABEL:fig:migrationpathsa) with fundamentally different characteristics. The first pathway, in-gap diffusion, involves Cr migration within the van der Waals (vdW) gap between Sb2Te3 quintuple layers, with low energy barriers and configurations similar to the training data, thereby testing interpolative accuracy.
@@ 3.3.2 Further Testing Different Trajectories
3.3.3 In-Gap Diffusion: A Test of Interpolative Accuracy
For the in-gap diffusion pathway (Fig. 2c–d), where the Cr atom moves within the van der Waals gap or just shallowly penetrates the interface, the atomic environments along the minimum energy path (MEP) are reasonably similar to the near-equilibrium states sampled during the AIMD simulations. In this regime, the performance hierarchy is clear. The fine-tuned models, particularly MACE FT–600K, demonstrate excellent agreement with the DFT reference, accurately predicting both the thermodynamic endpoints and the kinetic barrier (Fig. 2d). This success highlights that, when the task lies within the domain of the training data, fine-tuning effectively “sharpens” the generalist foundation model’s potential energy surface (PES) to capture system-specific details. In contrast, the MACE Scratch model, despite being trained on the same data, fails significantly, underscoring its inability to learn the subtle energy differences required to resolve the transition state from a limited dataset.
3.3.4 Deep Penetration: A Test of Extrapolative Robustness
@@ 3.4 Interlayer Sliding: Probing Robustness to Non-Local, Collective Displacements
Having tested the models’ response to local, high-distortion events, we evaluated their robustness to a different class of out-of-distribution challenge: large-scale, collective lattice displacements. To this end, we simulated the shearing of one Sb2Te3 layer relative to the other, a process governed by the subtle corrugations of the interlayer van der Waals potential energy surface (Fig. 3a). This task is particularly challenging for MLFFs as the local atomic environment of any given atom changes only minimally, while the global configuration undergoes a significant transformation that crosses periodic boundaries. The configurations along this sliding path were not present in the training data.
@@ 3.5 Representation Learning Analysis
To elucidate the origins of the observed performance differences, we conducted an analysis of the learned representations by projecting the high-dimensional atomic environment descriptors from each model into a two-dimensional space using t-distributed stochastic neighbor embedding (t-SNE), as shown in Fig. 4. This visualization reveals striking differences in how models trained with different strategies encode the same physical system.
The MACE Foundation and MACE Scratch models occupy distinct, well-separated regions of the latent space, as illustrated in the cross-model overlay (Fig. 4c) and quantified by high average silhouette scores (Fig. 4d). The foundation model’s representations are relatively compact, reflecting a generalist encoding trained to be robust across a vast chemical space. In contrast, the scratch-trained model produces a more dispersed representation, suggesting it has learned highly specific features to differentiate configurations within our limited in-house dataset; this specialization may contribute to its poor generalization on out-of-distribution tasks, such as NEB pathways.
Crucially, the representations of the two fine-tuned models (FT–600K and FT–Multi-T) are not only similar to each other (Fig. 4a, b) but also occupy a region of the latent space intermediate between the foundation and scratch models (Fig. 4c). This provides a powerful insight into the mechanism of fine-tuning: it effectively pulls the generalist representations of the foundation model towards the specialized representations required for the new task. Since fine-tuning begins with the foundation model’s weights and proceeds with the same in-house data used to train the scratch model, the resulting learned representations manifest as a hybrid.
This “hybrid representation” explains the balanced performance of the fine-tuned models. They retain sufficient general knowledge from the foundation model to remain stable and perform well on non-local tasks, while also incorporating specialized features from the in-house dataset that enable accurate prediction of system-specific properties, such as the in-gap migration barrier—outperforming both of the baseline models on that task. This finding underscores that model performance differences arise not from minor parameter adjustments, but from qualitatively distinct learned representations of the underlying material physics.
Figure 4: Representation learning analysis reveals distinct atomic environment encodings across training strategies. Two-dimensional t-SNE projections of atomic environment descriptors extracted from 600 K molecular dynamics trajectories. (a) The latent space embedding for the FT-600K model. (b) The embedding for the FT-Multi-T model, showing a similar distribution to the FT-600K model. (c) A cross-model overlay of representations from all four strategies. The foundation (red) and scratch-trained (orange) models occupy distinct, well-separated regions of the latent space. The fine-tuned models (dark and light blue) occupy an intermediate region, demonstrating that fine-tuning adapts the generalist representations towards the specialist domain. (d) Average silhouette scores for each model’s cluster relative to the others.
@@ 4 Conclusion
We benchmarked specialist (from-scratch) and generalist (foundation-based) machine-learned force fields (MLFFs) for Cr-intercalated Sb2Te3, emphasizing migration pathways as a diagnostic probe. NEB trajectories revealed that, while all models reproduce equilibrium structures, their performance diverges sharply for kinetic tasks. Migration thus provides a generalizable benchmark that tests both interpolation and extrapolation, exposing weaknesses invisible to equilibrium-only validation.
@@
These findings reframe the specialist vs. generalist debate: the choice is not just about efficiency but about qualitatively different physics learned by each model. Migration-based probes, coupled with latent-space diagnostics, offer a practical route to uncover hidden failures and guide uncertainty-aware active learning. This framework has the potential to generalize beyond the present case, enabling more robust and data-efficient MLFF development for accelerated materials discovery.

### v2

# AI $\workshoptitleAI for Accelerated Materials Discovery (AI4Mat) Workshop$

@@ Abstract
Machine-learned force fields (MLFFs), particularly pre-trained foundation models, are revolutionizing computational materials science by enabling ab initio-level accuracy at the length- and time-scales of classical molecular dynamics (MD). However, their rapid proliferation presents a critical strategic question: Should researchers train bespoke “specialist” models from scratch, fine-tune large “generalist” foundation models, or employ hybrid approaches? The trade-offs in data efficiency, predictive accuracy, computational cost, and susceptibility to out-of-distribution failure remain poorly understood, as does the fundamental question of how different training paradigms affect learned physical representations.
@@
Our key findings reveal that while all models adequately capture equilibrium structures, their predictions for non-equilibrium processes diverge dramatically. Targeted fine-tuning substantially outperforms both from-scratch and zero-shot approaches for kinetic properties, but induces catastrophic forgetting of long-range physics. Critically, analysis of learned representations shows that different training paradigms produce fundamentally distinct, non-overlapping latent space encodings, suggesting they capture different aspects of the underlying physics.
This work provides practical guidelines for MLFF development and establishes migration-based probes as an efficient, broadly applicable strategy for distinguishing model quality. We hope that this approach will offer a diagnostic framework that links performance to learned representations, paving the way for more intelligent, uncertainty-aware active learning strategies.
@@ 1 Introduction
The discovery and design of novel two-dimensional (2D) van der Waals (vdW) materials (Novoselov et al., 2016; Manzeli et al., 2017; Guo et al., 2021; Liu et al., 2024) continues to drive innovation in spintronics (He et al., 2022), quantum devices (Qian, 2024; Song and Gabor, 2018; Nayak, 2025), and energy technologies (Jin et al., 2019; Gautam et al., 2024). A particularly powerful strategy for tuning the properties of these layered materials involves doping—the insertion of guest atoms between vdW layers (Novoselov et al., 2016; Whittingham, 2004). This approach has enabled remarkable advances, from lithium-ion batteries (Winter et al., 1998; Hu et al., 2024) to the engineering of magnetic semiconductors (Gibertini et al., 2019) and quantum anomalous Hall systems (Tokura et al., 2019).
@@
Machine-Learned Force Fields (MLFFs) (Unke et al., 2021; Smith et al., 2017; Chen et al., 2017; Batatia et al., 2022; Vandermause et al., 2022; Batzner et al., 2022; Bartók et al., 2010; Zeng et al., 2023; Wines and Choudhary, 2025; Musaelian et al., 2023; Chen and Ong, 2022) have emerged as a powerful solution to this accuracy-versus-cost dilemma. The recent development of large-scale, pre-trained "foundation models" such as CHGNet (Deng et al., 2023), M3GNet/MEGNet (Chen et al., 2019), MACE-MP (Batatia et al., 2022), MatterSim (Yang et al., 2024), and UMA (Wood et al., 2025) marks a paradigm shift (Merchant et al., 2023). This shift presents researchers with a critical and unresolved dilemma: for a specific system, is it more effective to invest significant resources to build a robust specialist model from scratch (Zhang, 2018; Smith et al., 2019; Chmiela et al., 2019), or to adapt a generalist foundation model through fine-tuning? While generic benchmarks like Matbench (Dunn et al., 2020) validate their broad utility, their reliability for specific, high-fidelity applications is an open question (Friederich et al., 2021; Deringer et al., 2019). A valid concern is that these models may fail to capture nuanced interactions governing critical processes (Schran et al., 2021; Kapil et al., 2022; Grisafi et al., 2019). Fine-tuning is a promising solution, but naïve strategies can lead to catastrophic forgetting, where general knowledge is erased. Key questions of data efficiency (Janet et al., 2019; Lookman et al., 2019), stability (Zuo et al., 2020; Jinnouchi et al., 2019), and transferability (Zeni et al., 2021; Vandermause et al., 2020) for non-equilibrium processes remain (Bernstein et al., 2019; Zhang et al., 2019).
Furthermore, when an MLFF-driven simulation fails, the underlying cause is often treated as a "black box." The default response—indiscriminately adding more data—is a brute-force approach that wastes resources. We propose that the key to overcoming this is to diagnose failure using physical probes. Specifically, we argue that migration pathways, evaluated via Nudged Elastic Band (NEB) trajectories (Miskin et al., 2025; Ruttinger et al., 2022), provide a powerful and generalizable probe that simultaneously tests interpolation and extrapolation performance. Such migration-based probes impose stricter requirements than equilibrium-only tests, offering a sharper distinction between models and a path toward more intelligent, uncertainty-aware active learning.
@@
  1. 1.
[…]
  2. 2.
    What are the performance and stability trade-offs for fine-tuning strategies?
  3. 3.
    Can migration pathways serve as a generalizable and efficient probe to benchmark specialist versus generalist models, and can latent-space analysis reveal signatures of failure to guide data acquisition?
@@ 2.1 Data Generation and Model Training
[…]
All machine learning force field models were trained using the MACE architecture, (Batatia et al., 2022) which is an equivariant message-passing neural network designed for accurate representation of atomic interactions and many-body correlations.
@@
    FT-Multi_T: Foundation model fine-tuned on multi-temperature data
@@ 2.2 Evaluation Framework
Models were assessed through: (i) MD simulations at 300K and 600K to evaluate structural stability and transport properties, (ii) nudged elastic band calculations for migration barriers, and (iii) representation learning analysis using t-SNE visualization of physically interpretable descriptors. Performance metrics included force/energy accuracy, dynamic stability, barrier prediction errors, and latent space clustering quality. Detailed computational parameters are provided in Appendix B.
Figure 1: Conceptual Diagram. A flowchart illustrating two competing workflows: (a) training a specialist MLFF from scratch, and (b) fine-tuning a generalist foundation model. The diagram highlights the benchmarking questions this paper addresses.
@@ 3 Results and Discussions
3.1 Equilibrium Properties: Foundation for Comparison
We considered four candidate models: a “zero-shot” foundation model, a bespoke model trained from scratch, and two fine-tuned variants of the foundation model. To make consistent evaluations to benchmark the performance of these models, we first assessed the performance of each training strategy via their performance in MD simulations (Fig. 2a). This provides a strong test of each model’s ability to generate not only thermodynamic and structural properties but also accurately reproduce key dynamic, and transport properties. Each model was used to drive a 200 $ps$ simulation, with the resulting trajectories then subjected to the same set of post-processing analyses to evaluate their resulting key physical properties; shown in Figure 2.
Figure 2: Comprehensive Benchmarking of MLFF Performance in Molecular Dynamics Simulations. (a) A schematic of the unified evaluation protocol. Four candidate models—a zero-shot foundation model, a bespoke model trained from scratch, and two fine-tuned variants—are each used to drive a 200 ps MD simulation. The resulting trajectories are then subjected to a uniform set of post-processing analyses to evaluate key physical properties. (b) Thermodynamic stability, demonstrated by the evolution of temperature and pressure over the simulation, which remain stable around their target values for all models. (c) Thermal transport properties, showing the Heat Flux Autocorrelation Function (HFACF) and its running integral to compute thermal conductivity ($\kappa$). (d) The Mean Squared Displacement (MSD), used to assess atomic mobility and calculate the diffusion coefficient. (e) The Velocity Autocorrelation Function (VACF), which describes the system’s underlying dynamics.
All MACE models—foundation, scratch, and fine-tuned variants—successfully reproduce the equilibrium structure of Cr-doped Sb2Te3, with RDF analysis showing excellent agreement with AIMD references for all atomic pair correlations (Fig. 7, see Appendix A for detailed analysis). While all models maintain stable thermodynamic ensembles at 600K (Fig. 2c), the zero-shot foundation model exhibits a persistent pressure offset, likely due to its training on 0K equilibrium structures rather than finite-temperature configurations.
Despite similar performance on local structural (RDF) and short-time dynamic (VACF) properties (Fig. 2e), the models diverge significantly in their predictions of long-timescale transport phenomena. Fine-tuned models predict higher diffusion coefficients than both foundation and scratch models (Fig. 2d), with the multi-temperature fine-tuned variant showing the largest enhancement. This suggests that exposure to high-temperature training data creates a flatter potential energy (PES) that persists even at lower temperatures.
@@ 3.1 Equilibrium Properties: Foundation for Comparison
3.2 Non-Equilibrium Diffusion Pathways: The Critical Test of Generalization
@@
3.2.1 Local Migration Events
@@
The Critical Role of Task-Specific Fine-Tuning
@@
To probe the generalization capabilities of the different MLFF training strategies, we evaluated their performance on two distinct Cr migration pathways (Fig. 3a) with fundamentally different characteristics. The first pathway, in-gap diffusion, involves Cr migration within the vdW gap between Sb2Te3 quintuple layers, with low energy barriers and configurations similar to the training data, thereby testing interpolative accuracy.
@@
In-Gap Diffusion: A Test of Interpolative Accuracy
For the in-gap diffusion pathway (Fig. 3c–d), where the Cr atom moves within the vdW gap or just shallowly penetrates the interface, the atomic environments along the minimum energy path (MEP) are reasonably similar to the near-equilibrium states sampled during the AIMD simulations. In this regime, the performance hierarchy is clear. The fine-tuned models, particularly MACE FT–600K, demonstrate excellent agreement with the DFT reference, accurately predicting both the thermodynamic endpoints and the kinetic barrier (Fig. 3d). This success highlights that, when the task lies within the domain of the training data, fine-tuning effectively “sharpens” the generalist foundation model’s potential energy surface (PES) to capture system-specific details. In contrast, the MACE Scratch model, despite being trained on the same data, fails significantly, underscoring its inability to learn the subtle energy differences required to resolve the transition state from a limited dataset.
Deep Penetration: A Test of Extrapolative Robustness
@@ 3.3 Interlayer Sliding: Probing Robustness to Non-Local, Collective Displacements
Having tested the models’ response to local, high-distortion events, we evaluated their robustness to a different class of out-of-distribution challenge: large-scale, collective lattice displacements. To this end, we simulated the shearing of one Sb2Te3 layer relative to the other, a process governed by the subtle corrugations of the interlayer vdW potential energy surface (Fig. 4a). This task is particularly challenging for MLFFs as the local atomic environment of any given atom changes only minimally, while the global configuration undergoes a significant transformation that crosses periodic boundaries. The configurations along this sliding path were not present in the training data.
@@ 3.4 Representation Learning Analysis
To elucidate the origins of the observed performance differences, we conducted a multi-faceted analysis of the learned representations. We projected the high-dimensional atomic environment descriptors from each model’s 600 K MD trajectory into a two-dimensional space using two complementary techniques: t-distributed stochastic neighbor embedding (t-SNE) (Maaten and Hinton, 2008) to assess representational dissimilarity, and Potential of Heat-diffusion for Affinity-based Trajectory Embedding (PHATE) (Moon et al., 2019) to reveal the underlying geometric structure of the system’s dynamics (Fig. 5).
The t-SNE projection (Fig. 5a) confirms that models trained with different strategies learn qualitatively distinct encodings. The MACE Foundation and MACE Scratch models occupy well-separated regions of the latent space, a finding quantitatively supported by their high average silhouette scores (Fig. 5d). This highlights the fundamental dissimilarity between the generalist prior and the specialist model trained from scratch. Crucially, the two fine-tuned models occupy an intermediate region, demonstrating that fine-tuning acts as a bridge between these two representational paradigms.
While t-SNE shows that the representations are different, PHATE reveals why it matters for physical prediction. The PHATE projection (Fig. 5b) maps the continuous time-evolution of the system to a low-dimensional "dynamical manifold"—an arc-like structure whose geometry is intrinsically linked to the potential energy surface (Fig. 5c). Along the primary axis of this manifold (PHATE 1), all specialist models (Scratch and FT) trace a similar region, indicating they have all learned the dominant, low-energy dynamics of the system.
The key insight comes from the separation along the second axis (PHATE 2). The scratch-trained model’s representation is isolated from fine-tuned models in a distinct region (higher PHATE 2 values). This suggests it has learned a brittle, overfit representation, effectively memorizing a narrow path through the energy landscape without understanding the broader physical context. In stark contrast, the fine-tuned models are constrained to a different region of the manifold. This is the signature of regularization by pre-training; the foundation model’s robust physical prior prevents the fine-tuned models from collapsing into the brittle state of the scratch model. Instead, they learn the system-specific dynamics within the context of a general and smooth physical manifold.
This geometric difference in the latent space provides a direct mechanistic explanation for the models’ performance on the diffusion task. Diffusion is a rare-event process that requires the model to generalize to out-of-distribution transition states. The scratch model’s isolated manifold corresponds to a noisy and untrustworthy PES outside its training domain, leading to unphysical dynamics and a flattened MSD plot. Conversely, the fine-tuned models’ robust, regularized manifold corresponds to a globally smoother and more reliable PES. This enables them to accurately predict the energy barriers and forces along the diffusion pathway, resulting in physically correct, linear MSD behavior. Fine-tuning succeeds not by creating a simple hybrid, but by aligning a general physical prior with the specific dynamical manifold of a target system, thereby enabling robust generalization to complex, long-timescale phenomena.
Figure 5: Representation Analysis Reveals How Fine-Tuning Aligns a Physical Manifold to Enable Generalization. Projections of atomic environment descriptors from 600 K MD trajectories. (a) t-SNE projection shows that Foundation (red) and Scratch (orange) models produce highly separated representations, while fine-tuned models (blue) are intermediate. (b) PHATE projection reveals the system’s continuous dynamical manifold. Note the clear separation between the brittle scratch model representation and the regularized fine-tuned models along the vertical axis (PHATE 2). (c) The same PHATE embedding colored by potential energy confirms that the manifold’s geometry corresponds to the physical energy landscape. The foundation model’s large energy offset indicates its nature as an uncalibrated prior. (d) Average silhouette scores (from t-SNE) quantify the representational dissimilarity. The results demonstrate that fine-tuning succeeds by constraining a generalist representation to the specific physical manifold of the target system, a process that regularizes the model and enables accurate prediction of complex dynamics like diffusion.
@@ 4 Conclusion
We benchmarked specialist (from-scratch) and generalist (foundation-based) machine-learned force fields (MLFFs) for Cr-doped Sb2Te3, emphasizing migration pathways as a diagnostic probe. NEB trajectories revealed that, while all models reproduce equilibrium structures, their performance diverges sharply for kinetic tasks. Migration thus provides a generalizable benchmark that tests both interpolation and extrapolation, exposing weaknesses invisible to equilibrium-only validation.
@@
These findings reframe the specialist vs. generalist debate: the choice is not just about efficiency but about qualitatively different physics learned by each model. Foundation models excel at capturing broad chemical trends but require careful adaptation for system-specific kinetics, while specialist models offer more reliable local accuracy at the cost of transferability. Migration-based probes, coupled with latent-space diagnostics, offer a practical route to uncover hidden failures and guide uncertainty-aware active learning.
Limitations and Future Work
Our current study focuses on a single material system and migration mechanism as a proof-of-concept case study. Future work should extend this framework to multiple material classes, investigate uncertainty quantification methods for identifying unreliable predictions, and develop hybrid approaches that combine the transferability of foundation models with the precision of specialist training. Additionally, exploring active learning strategies guided by migration-based diagnostics could enable more efficient data collection for kinetic properties. This framework has the potential to generalize beyond the present case, enabling more robust and uncertainty-aware MLFF development for accelerated materials discovery.